OpenStudy (90890908967777878888):

express 2.6 as mixed number in lowest terms

OpenStudy (alexandervonhumboldt2):

2.6=26/10=(2*13)/(2*5): quit 2: 13/5=2 3/5 a b/c=(a*c+b)/c

OpenStudy (90890908967777878888):

what the thhe answer